# Multi-Agent Target Assignment and Path Finding for Intelligent Warehouse: A Cooperative Multi-Agent Deep Reinforcement Learning Perspective

### v1

@@ 5.2 Cooperation Ability
We also designed a conflict scenario to verify the learned cooperation ability of agents. As shown in Figure 11, the big purple circle represents agent-1, its task is denoted by the small purple circle. The big gray circle represents agent-2, its task is denoted by the small black circle. We deliberately block other roads to create the conflict environment. As shown in Figure 11, there is bound to be a conflict between agent-1 and agent-2 during the navigation. The trajectories generated by our method are shown in Figure 11, the red curve is the trajectory of agent-1, and the cyan curve is the trajectory of agent-2. Results show that both agent-1 and agent-2 had learned to avoid each other at the point of conflict, and then navigated to their tasks. This verified the cooperation ability of our method.
@@ 5.3 Time Efficiency
Table 1 shows the time comparisons between traditional methods [35, 36] and our method. Table 1 shows that our method simultaneously addresses the target assignment problem (TA) and the path planning (PF), however, the traditional method addresses the target assignment first and then performs the path planning. Thus, the consumed time in our method is TAPF time; however, the consumed time in the traditional method is TA+PF. In Table 1, ES means each step. The reason that we compare the consumed time of each step is: (1) Our method handles continuous action space problem; however, traditional method handles discrete action space problem. (2) The consumed time of path finding can be influenced by many factors, such as the resolution of the grid map and the distance taken at each step in the grid map. (3) previous literature [35, 36] did not consider the physical dynamics of agents. However, the physical dynamics of agents has been considered in this study. Thus, for a fair comparison, we only compare the consumed time of each step, which means the consumed time from an agent obtain an observation to output the decision action. For traditional method, the consumed time of target assignment is computed by the method in [35], and the consumed time of each step in path planning is computed by the method in [36].
Table 1 shows that in easy tasks such two agents - two tasks and two agents - four tasks, the consumed time of target assignment in traditional method can be acceptable. However, with the difficulty of tasks increased, the time of solving target assignment in traditional method increased rapidly, especially in the five agents - twenty tasks scenario. However, our method provides policy time-efficiently in all different level tasks (from two agents - two tasks to five agents - twenty tasks). This verified the time-efficiency ability of our method. As we all know, time efficiency is an important factor in real-world TAPF problems. Therefore, the results of the experiment demonstrated the promising engineering practical value of our method. Note that Figure 6 to Figure 10 had shown the performances of target assignment and path finding of our method in five different level scenarios. For target assignment, the results show that task assignment is reasonably addressed because our method assigns the task to the agent that is close to the task. For path finding, the results show that the planned paths are almost shortest.
@@ 6 Conclusion and future work
In this study, we propose a method to simultaneously address the TAPF problem in intelligent warehouse from a perspective of cooperative multi-agent deep RL. First, we model the TAPF problem as cooperative multi-agent deep RL problem. Then, we simultaneously address target assignment and path finding by cooperative multi-agent deep RL algorithm. Moreover, previous literature rarely considers the physical dynamics of agents. In this study, the physical dynamics of agents have been considered. Experimental results show that our method performs well in various of task settings, which means the target assignment is solved reasonably and the planned path is almost shortest. Furthermore, our method is more time-efficient than baselines. For future work, we will apply our method to real-world target assignment and path finding problem.
7 Acknowledgments
This research was funded by Shenzhen basic research program JCYJ20180507183837726 and National Natural Science Foundation U1813206, 61977019.
@@ 7 Acknowledgments
Network Architecture: The actor network contains four fully connected layers. As described in Section 4.1, the input dimension of the first layer is equal to the observation dimension. The hidden dimension is 128. The first three layers are followed by ReLU activation function. The output dimension of the forth layer is the action dimension. The critic network contains four fully connected layers. The input dimension of the first layer is equal to the observation dimension plus the action dimension. The hidden dimension is 128. The first three layers are followed by ReLU activation function. The output of the forth layer is the state-action value. Networks are trained by Adam optimizer.

### v2

@@ 5.2 Cooperation Ability
We also designed a conflict scenario to verify the learned cooperation ability of agents. As shown in Figure 11, the big purple circle represents agent-1, its task is denoted by the small purple circle. The big gray circle represents agent-2, its task is denoted by the small black circle. We deliberately block other roads to create a conflict environment. As shown in Figure 11, there is bound to be a conflict between agent-1 and agent-2 during the navigation. The trajectories generated by our method are shown in Figure 11, the red curve is the trajectory of agent-1, and the cyan curve is the trajectory of agent-2. Results show that both agent-1 and agent-2 had learned to avoid each other at the point of conflict, and then navigated to their tasks. This verified the cooperation ability of our method.
@@ 5.3 Time Efficiency
Table 1 shows the time comparisons between traditional methods [35, 36] and our method. Table 1 shows that our method simultaneously addresses target assignment (TA) and path planning (PF) problems, however, the traditional method addresses target assignment first and then performs path planning. Thus, the consumed time in our method is TAPF time, however, the consumed time in the traditional method is TA+PF. In Table 1, ES means each step. The reasons that we compare the consumed time of each step is: (1) Our method handles continuous action space problems, however, the traditional method handles discrete action space problems. (2) The consumed time of path finding can be influenced by many factors, such as the resolution of the grid map and the distance taken at each step in the grid map. (3) previous literature [35, 36] did not consider the physical dynamics of agents. However, the physical dynamics of agents have been considered in this study. Thus, for a fair comparison, we only compare the consumed time of each step, which means the consumed time from an agent obtains an observation to output the decision action. For the traditional method, the consumed time of target assignment is computed by the method in [35], and the consumed time of each step in path planning is computed by the method in [36].
Table 1 shows that in easy tasks such two agents - two tasks and two agents - four tasks, the consumed time of target assignment in the traditional method can be acceptable. However, as the difficulty of tasks increased, the time of solving target assignments in the traditional method increased rapidly, especially in the five agents - twenty tasks scenario. However, our method provides policy time-efficiently in all different level tasks (from two agents - two tasks to five agents - twenty tasks). This verified the time efficiency ability of our method. As we all know, time efficiency is an important factor in real-world TAPF problems. Therefore, the experiment results demonstrated the promising engineering practical value of our method. Note that Figure 6 to Figure 10 had shown the performances of target assignment and path finding of our method in five different level scenarios. For target assignment, results show that the task assignment is addressed reasonably because our method assigns the task to the agent that is close to the task. For pathfinding, results show that the planned paths are almost shortest.
@@ 6 Conclusion and future work
In this study, we propose a method to simultaneously address the TAPF problem in the intelligent warehouse from a perspective of cooperative multi-agent deep RL. First, we model the TAPF problem as a cooperative multi-agent deep RL problem. Then, we simultaneously address target assignment and pathfinding by cooperative multi-agent deep RL algorithm. Moreover, previous literature rarely considers the physical dynamics of agents. In this study, the physical dynamics of agents have been considered. Experimental results show that our method performs well in various task settings, which means the target assignment is solved reasonably and the planned path is almost the shortest. Furthermore, our method is more time-efficient than baselines. For future work, we will apply our method to real-world target assignments and pathfinding problems.
@@
Network Architecture: The actor network contains four fully connected layers. As described in Section 4.1, the input dimension of the first layer is equal to the observation dimension. The hidden dimension is 128. The first three layers are followed by ReLU activation function. The output dimension of the fourth layer is the action dimension. The critic network contains four fully connected layers. The input dimension of the first layer is equal to the observation dimension plus the action dimension. The hidden dimension is 128. The first three layers are followed by ReLU activation function. The output of the fourth layer is the state action value. Networks are trained by Adam optimizer.